OpenStudy (anonymous):

@JamesJ CALC 2!!! Find the MacLaurin Polynomial of degree 5 for f(x)= the integral from 0 to x of (e^((-5t)^4)) dt

OpenStudy (anonymous):

please help im lost.

OpenStudy (jamesj):

e^((-5t)^4 = e^(-20t) Integrate that function from 0 to x. Remember also that the MacLaurin poly of e^ax is \[e^{ax} = 1 + ax + (ax)^2/2! + (ax)^3/3! + (ax)^4/4! + (ax)^5/5! + ... \]

OpenStudy (turingtest):

Hey James can you help really quick on something that seems very simple, but is driving me crazy? it should only take a moment.